The following cases makes different assumptions with respect to the amounts of income and deductions of Mr. Michel Daigle for the current taxation year under the various subdivisions of Division B of the Income Tax, and the last two previous years’ taxes payable:

Case 2-A

Employment Income

$35,000

Income (Loss) From Business

(10,000)

Income From Property

12,000

Taxable Capital Gains

42,000

Allowable Capital Losses

(18,000)

Subdivision E Deductions (Spousal Support)

(4,000)

Taxable Income Previous Years:

2016: $10,000

2017: nil

Case 2-B

Employment Income

$33,000

Income (Loss) From Business

(39,000)

Income From Property

14,000

Taxable Capital Gains

36,000

Allowable Capital Losses

(42,000)

Subdivision E Deductions (Spousal Support)

(7,000)

Taxable Income Previous Years:

2016: $10,000

2017: $8,000

Required: For Both cases, calculate Mr. Daigle’s Net Income for tax purposes, indicate any possible loss carryovers from the current year, taxes payable for current year, and calculate require installment payment for current year.

(assume Federal rate of 15% and Provincial Rate of 7%)
Answer #1

Answer

CASE 2-A

Given

Employment Income=35000

Income loss from business=(10000)

Income from houseproperty=12000

Taxable capital gain 42000

Allowable capital losses=(18000)

Subdivision E deduction=(4000)

* capital loss is a allowable expense from capital gain and hence deductable.

Mr diagles net income for tax purposes:-

Enploymemt income 35000
Less Income loss from business (10000)
Add Income from property 12000
Add Taxable capital gain 42000
Less Allowable capital loss (18000)
Less Subdivision E deduction (4000)
Total income for tax purposes:- 57000

Possible loss carryover

In case 2-A there is no possible loss carryover ,all are set off against income.

Tax payable case 2-A

federal rate=15%=57000×15%=8550

Provincial tax=7%=57000×7%=3990

Tax payable= 8550+3990=12540

Case 2-B

Employment income 33000
Income from property 14000
subdivision E deduction (7000)
Total income for tax purposes:- 40000

*Capital gain loss carryover=(42000)-36000=(6000)

*loss from bussiness can not be set off from employment income

*loss of business income of 39000 can be set off from previous year taxable income=39000-10000-8000=21000( total loss carryover next year)

*Tax payable

federal tax=15%=40000×15%=6000

Provincial tax=40000×7%=2800